This is the first problem x+4>5 or x+4<-2 The next problem is Find three consecutive integers such that the sum of the first and third is -38 I am so thankful that I have found you. You are a blessing
Answer by ankor@dixie-net.com(22740)


x + 4 > 5
:
Subtract 4 from both sides;
x > 5 - 4
x > 1
:
or x + 4 <- 2
Subtract 4 from both sides
x <= - 2 - 4
x <= -6
:
:
Find three consecutive integers such that the sum of the first and third
is -38.
Let the 3 consecutive intergers = x, x+1, x+2
:
Sum of the 1st (x) and the third (x+2) = -38
x + (x+2) = -38
2x + 2 = -38
:
Subtract 2 from both sides:
2x = -38 - 2
2x = -40/2
x = -20 is the 1st digit, then -19, -18
:
Check: Sum of 1st and 3rd digits: -20 + -18 = -38